# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

### v1

@@ Abstract.
Multi-agent systems - systems with multiple independent AI agents working together to achieve a common goal - are becoming increasingly prevalent in daily life. Drawing inspiration from the phenomenon of human group social influence, we investigate whether a group of AI agents can create social pressure on users to agree with them, potentially changing their stance on a topic. We conducted a study in which participants discussed social issues with either a single or multiple AI agents, and where the agents either agreed or disagreed with the user’s stance on the topic. We found that conversing with multiple agents (holding conversation content constant) increased the social pressure felt by participants, and caused a greater shift in opinion towards the agents’ stances on each topic. Our study shows the potential advantages of multi-agent systems over single-agent platforms in causing opinion change. We discuss design implications for possible multi-agent systems that promote social good, as well as the potential for malicious actors to use these systems to manipulate public opinion.
Social Influence, Multi-agent, LLM Agent, Persuasion
*Equal contribution
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 18–22, 2025; XXX, XXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Psychologyccs: Human-centered computing User studies
Figure 1. Overview of our study.
@@ 1. Introduction
AI agents are advanced systems capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). These agents leverage the conversational and reasoning capabilities of large language models (LLMs) like Gemini, GPT-4, and LLaMA, and have become commonplace in daily life (Reuters, 2024). Significant research has also explored using these agents across a broad range of scenarios including education (Lieb and Goel, 2024), search interfaces (Ma et al., 2024), and trading platforms (Yu et al., 2024). The prevalence and ubiquity of LLM-powered agents has given rise to the concept of multi-agent systems - systems with two or more independent agents working together to fulfill a common goal and improve task performance (Guo et al., 2024; Chen et al., 2023b; Clarke et al., 2022). Major AI companies such as Anthropic (Claude, 2024) and Google (Google, 2024) have recently introduced multi-agent frameworks, making it easier for developers to create such interfaces.
@@
With both the rise of multi-agent systems and the increasing prevalence of virtual agents in our daily lives, an interesting question emerges: Can a group of AI agents collectively exert social influence on users? This arises from a well-documented observation in human-human interaction: individuals are more likely to be influenced by the opinions of larger groups of people, a phenomenon known as social influence. Various forms of social influence, including conformance to the majority (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), desire for social acceptance (Richters and Waters, 1991; Oostveen et al., 1996), and adherence to social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011), are pervasive in everyday life. It is well-known that human-computer interaction dynamics tend to mimic human-human dynamics (Nass et al., 1994), with humans perceiving and reacting to computer systems similar to how they would behave towards other humans. This leads to the possibility that a group of AI agents holding the same opinion may exert social influence on users, similar to how a group of humans would.
@@
The importance of investigating this issue is two-fold. Firstly, multi-agent interactions have become increasingly common. Most notably, on social media platforms like Twitter, human-like social bots have been effectively used to disseminate information and increase user engagement (Wischnewski et al., 2024). Multi-agent systems have also been designed to support online discussions (Zhang et al., 2024), public information elicitation (Jiang et al., 2023), and health coaching (Beinema et al., 2021), exposing users to opinions from multiple agents in a range of settings. Secondly, the potential social influence of multi-agent systems on users raises potential concerns. Evidence from psychology studies suggests that interactions with groups sharing similar views can significantly impact user opinions (Myers and Lamm, 1976; Isenberg, 1986), potentially leading to opinion polarization and manipulation. As digital agents increasingly function as social actors, the influence they exert on users is a critical yet underexplored issue.
@@
Beyond potential negative effects, the social influence of multi-agent systems can also be harnessed in a positive way. In the traditional human-human domain, majority opinion has long been used to encourage changes in attitudes and behaviors, such as in health interventions (Skalski and Tamborini, 2007; Zhang et al., 2015) and sustainability efforts (Athanasiadis and Mitkas, 2005; Vossen et al., 2009). In human-computer interaction, persuasive systems have been designed across numerous domains like health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021), and social welfare (Shi et al., 2020), aiming to improve user behavior or change their attitudes through persuasion and informational education. Yet, these implementations are limited to single-agent interactions. The potential of multi-agent systems that collectively effect changes in user behavior or opinion remains unexplored.
@@
The potential of AI agents in enacting persuasive effects is further supported by research in the HCI domain, which has highlighted the significant social influence individual AI agents exert, shaping users’ attitudes (Jakesch et al., 2023) and behaviors (Shi et al., 2020; Zhang et al., 2020). However, such research has focused on the design characteristics of individual agents: for instance, on how the perceived identity of an agent as a chatbot or a human impacts outcomes (Shi et al., 2020). Little attention has been given to how the number of agents affects the effectiveness of these systems. Existing research on multi-agent system has largely revolved around enhancing task performance (Guo et al., 2024; Chen et al., 2023b), either by helping users complete tasks more effectively or improving the technical performance of the system itself (Du et al., 2023; Chen et al., 2023a). In contrast, the indirect effects of social influence caused by these multi-agent systems has been largely overlooked. The impact of social influence is less relevant in these systems, where the primary goal is often to improve the clarity of information conveyed to the user. In scenarios where the system needs to both convey information and persuade a user of a certain viewpoint, the effect of multiple vs. single agents is critically underexplored.
@@
Therefore, given 1) the increasing prevalence of multi-agent systems and interactions with multiple social bots online, 2) the known social influence that AI agents can exert, and 3) the established impacts of group social influence among humans that may be mirrored in human-AI interactions, there is 4) significant potential for both positive and negative effects caused by the collective social influence created by a group of AI agents. In this study, we address this issue by asking: Can interacting with multiple virtual agents create social influence that shifts our opinions? Specifically, we explore whether interactions with multiple agents can induce greater levels of social influence and lead to stronger opinion changes, and identify demographics factors that increase susceptibility to AI influence, and aim to form guidelines for the effective and ethical implementation of multi-agent systems.
@@
To investigate these questions, we conducted an experiment (n = 94) where participants discussed their stances on two social topics with one, three, or five agents. Each participant was randomly assigned to a condition, with discussion content kept consistent across groups, differing only in the number of agents presenting it. Agents presented arguments that aligned with the participant’s stance on one topic and opposed it on the other. Survey data and open-ended responses were collected before and after each conversation round.
@@
Through quantitative and qualitative analysis, we found that the multi-agent setup significantly influenced opinion shifts. Participants showed stronger opinion changes toward the agents when they disagreed and greater polarization when agents agreed with them. However, the shift did not intensify as the number of agents increased from 3 to 5. Instead, the 5-agent group led to a sense of polarization, with more participants rejecting the agents’ arguments and moving further from the agents’ positions. Additionally, multi-agent groups heightened perceptions of social influence, particularly normative influence to align with the majority. Open-ended responses indicated that participants expressed a desire for affiliation within the group, which contributed to social influence and opinion shifts in the 3- and 5-agent conditions. Demographic analysis further revealed that younger participants were more susceptible to multi-agent influence.
@@
Our study makes the following contributions:
@@
  • We provide a foundational understanding of the differences in social influence between single-agent and multi-agent systems. By observing participants’ feelings of social exclusion and their desire to engage, we identify an implicit sense of affiliation with a group of agents, expanding existing social influence theories to include human-agent interactions.
  • We provide design suggestions for using multi-agent setups to more effectively convey persuasive strategies. By leveraging the normative influence exerted by multiple agents, future systems could foster emotional engagement in participants, promoting more lasting and internalized changes in attitudes and behaviors.
@@ 2. Related Work
2.1. Multi-Agent Interactions
AI agents are increasingly integrated into our daily life. These advanced AI systems are capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). Recent research and industry applications have leveraged the conversational and reasoning capabilities of LLMs to create problem-solving tools (Wang et al., 2024a; Roy et al., 2024; Ning et al., 2024; Yu et al., 2024; Guan et al., 2024), educational assistants (Lieb and Goel, 2024), and novel search interfaces (Ma et al., 2024). In the industry, many companies deploy AI agents on social media to assist with branding and customer service (Xu et al., 2017; Hu et al., 2018; Jiang et al., 2022), while others embed agents in their applications for assisting writing (Dhillon et al., 2024), brainstorming (Rezwana and Maher, 2023), or gaming (Sidji et al., 2024).
Despite this, there are many tasks that a single LLM-based agent struggles to handle effectively. Individual LLMs can lack sufficient domain expertise to handle specialized tasks (Ge et al., 2023), and may have insufficiently robust reasoning capabilities. Inspired by existing psychological theories (Smoliar, 1991), studies have investigated techniques such as ”debating” (Du et al., 2023) or ”reconciling” (Chen et al., 2023a) between multiple agents, to enhance reasoning by leveraging diverse perspectives. This has driven growing interest in ”multi-agent collaboration,” where agents work together to improve task performance (Guo et al., 2024; Chen et al., 2023b). Fields such as natural language processing (Guo et al., 2024), software engineering (Cardoso and Ferrando, 2021), and robotics (Chen et al., 2024) have leveraged the collective intelligence of multiple agents to simulate a group dynamic and improve task performance (Qian et al., 2023; Hong et al., 2023; Xiong et al., 2023). Multi-agent setups have also been used to simulate different human roles, akin to multiple workers interacting (Park et al., 2022, 2023b) or collaborating on a task (Light et al., 2023; Aher et al., 2023).
@@ 2.1. Multi-Agent Interactions
Given the potential prevalence of human and multi-agent collaboration in everyday life, the interaction between multiple agents and humans remains underexplored in the HCI domain. Previous work has a limited focus on multi-agent systems in HCI, mostly highlighting opportunities of multi-agent interfaces in improving user outcomes: either by reducing cognitive load in information gathering (Jiang et al., 2023) or to provide information from diverse sources to improve decision making (Tan and Liew, 2022; Park et al., 2023a; Beinema et al., 2021; Chaves and Gerosa, 2018). These studies also show the challenges that multi-agent interfaces create for users, such as causing users to feel overwhelmed or confused by multiple agents providing different inputs (Chaves and Gerosa, 2018; Tan and Liew, 2022; Park et al., 2023a). In contrast, no existing work to our knowledge has investigated if and how a group of agents affects users through social influence mechanisms. Such effects are distinct from the cognitive and decision making effects in previous studies, yet are potentially as important.
In this study, we investigate the scenario: what if multiple agents are perceived as social actors, and they all share the same opinion on a topic just like a group of people might? This situation can arise, for instance, when companies deploy multiple agents on social media with a unified stance. While the influence of a single opinionated bot is well-documented (Jakesch et al., 2023; Shi et al., 2020; Zhang et al., 2020), the impact of being surrounded by multiple like-minded bots is less understood. Considering the growing prevalence of multi-agent systems, it is crucial to investigate how they might shape people’s opinions. Drawing from existing theories on human behavior within social groups, we hypothesize several potential effects, which we discuss in the next section.
@@ 2.2. Social Influence Theory
When multiple people express the same opinion, others in the group often feel compelled to agree. This phenomenon is known as ”social influence” (Cialdini and Goldstein, 2004), which comprises the ways in which individuals adjust their behavior to meet the demands of a social environment. This theory is well-known and widely studied in social psychology, and has been applied across different contexts, such as marketing (Salganik et al., 2006; Zhu et al., 2012; Teo et al., 2019), health intervention (Skalski and Tamborini, 2007; Zhang et al., 2015), sustainability (Athanasiadis and Mitkas, 2005; Vossen et al., 2009) and political discussions (Price et al., 2006).
@@
Social influence is a broad term that manifests in various forms. Two main types of social influence are normative influence, the desire to ”conform with the expectations of another”, and informational influence, the desire to ”accept information obtained from another as evidence about reality” (Deutsch and Gerard, 1955). Normative influence relates to many common types of social influence including conformity (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), socialization (Richters and Waters, 1991; Oostveen et al., 1996), and social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011). More recent literature has termed this mechanism affiliation (Cialdini and Goldstein, 2004), or a desire to fit in with the majority: assimilation into a social group is often desired, and agreeing with the majority opinion can expedite this process. Naturally, normative social influence is more common among individuals that form a collective social group (Deutsch and Gerard, 1955), which creates an environment that one might want to be a part of. In this study, we investigate whether a group of AI agents can create a similar sense of a social group, implicitly encouraging a user to join the group by way of agreeing with their opinions.
Informational influence, on the other hand, relates to the concept of accuracy (Cialdini and Goldstein, 2004), or the natural instinct to gravitate towards objectively correct opinions as expressed by a group. This is distinct from normative influence as individuals may ”accept an opponent’s beliefs as evidence about reality even though one has no [other] motivation to agree with him” (Deutsch and Gerard, 1955). Despite this, social influence can still persuade individuals to accept an opinion, as the perceived consensus (Cialdini and Goldstein, 2004) of a larger number of people expressing the same viewpoint may suggest that that viewpoint is accurate. One distinction between human-human and human-AI interactions is that in human groups, each person represents a distinct source of knowledge and possible truth; with AI agents, it is trivial to set up multiple agents based on the same source of data, meaning that the number of agents presenting an opinion has no relation to its objective accuracy. Nevertheless, the tendency of humans to perceive such agents as distinct social actors (Nass et al., 1994) raises the possibility of informational influence from a group of agents.
@@
In HCI research, a handful of studies (e.g. (Adam et al., 2021)) have examined the social influence that individual chatbots can exert on humans. In this study, we aim to conduct, to our knowledge, the first investigation of the possibility of both normative and informational social influence caused by groups of AI agents.
2.3. Conversational Agents in Online Discussions

### v2

@@ Abstract.
Multi-agent systems, systems with multiple independent AI agents working together to achieve a common goal, are becoming increasingly prevalent in daily life. Drawing inspiration from the phenomenon of human group social influence, we investigate whether a group of AI agents can create social pressure on users to agree with them, potentially changing their stance on a topic. We conducted a study in which participants discussed social issues with either a single or multiple AI agents, and where the agents either agreed or disagreed with the user’s stance on the topic. We found that conversing with multiple agents increased the social pressure felt by participants, and caused a greater shift in opinion towards the agents’ stances on the conversation topics. Our study shows the potential advantages of multi-agent systems over single-agent platforms in causing opinion change. We discuss the resulting possibilities for multi-agent systems that promote social good, as well as potential malicious actors using these systems to manipulate public opinion.
Social Influence, Multi-agent, LLM Agent
*The two authors contributed equally to this work.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW452publicationmonth: 11doi: 10.1145/3757633ccs: Applied computing Psychologyccs: Human-centered computing User studies
@@ 1. Introduction
AI agents are advanced systems capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). These agents leverage the conversational and reasoning capabilities of large language models (LLMs) like Gemini, GPT-4, and LLaMA, and have become commonplace in daily life (Reuters, 2024). Significant research has also explored the use of such agents in a broad range of scenarios, including education (Lieb and Goel, 2024), search interfaces (Ma et al., 2024), and trading platforms (Yu et al., 2024). The prevalence and ubiquity of LLM-powered agents have given rise to the concept of multi-agent systems-systems with two or more independent agents working together to fulfill a common goal and improve task performance (Guo et al., 2024; Chen et al., 2023b). Major AI companies such as Anthropic (Claude, 2024) and Google (Google, 2024) have introduced multi-agent frameworks, making it easier for developers to create such interfaces.
@@
With both the rise of multi-agent systems and the increasing prevalence of virtual agents in our daily lives, an interesting question emerges: Can a group of AI agents collectively exert social influence on users? This question arises from a well-documented observation in human-human interaction: individuals are more likely to be influenced by the opinions of larger groups of people, a phenomenon known as social influence. Various forms of social influence, including conformity to the majority (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), desire for social acceptance (Richters and Waters, 1991; Oostveen et al., 1996), and adherence to social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011), are pervasive in everyday life.
While prior work in HCI has argued that human-computer interaction often mirrors human-human dynamics—where users perceive and respond to computers as social entities (Nass et al., 1994)—AI agents differ from humans in two important ways: First, regarding knowledge, AI agents are often perceived as having broader but more generic knowledge, particularly in factual or domain-specific areas, while lacking personal, contextual, or experiential insight (Luger and Sellen, 2016). Unlike human groups, where informational influence tends to increase with group size (Cialdini and Goldstein, 2004), it is unclear whether users perceive a group of AI agents as more informational than a single agent. Second, in terms of social characteristics, AI agents often lack key social attributes (Ozdemir et al., 2023) such as empathy, credibility, and authority, which are typically foundational to social influence in human groups (Cialdini and Goldstein, 2004). Consequently, it remains uncertain whether multiple AI agents can evoke the same sense of normative pressure or emotional affiliation that human groups do. Given these differences, it remains an open question whether multiple AI agents can exert social influence in ways comparable to human social groups. This leads to a central research question of our work: Can a group of AI agents holding the same opinion influence user decisions in a manner similar to human group influence?
@@
The importance of this question is twofold. Firstly, multi-agent interactions have become increasingly common, most notably on social media platforms like X (formerly Twitter), where human-like social bots have been effectively used to disseminate information and increase user engagement (Wischnewski et al., 2024). Multi-agent systems have also been designed to support online discussions (Zhang et al., 2024), public information elicitation (Jiang et al., 2023), and health coaching (Beinema et al., 2021), exposing users to opinions from multiple agents in a range of settings. Secondly, the potential social influence of multi-agent systems on users would raise concerns. Evidence from psychology studies suggests that interactions with groups sharing similar views can significantly impact user opinions (Myers and Lamm, 1976; Isenberg, 1986), potentially leading to opinion polarization and manipulation. As digital agents increasingly function as social actors, the influence they exert on users is a critical yet underexplored issue.
@@
Beyond potential negative effects, the social influence of multi-agent systems can also be harnessed in a positive way. The social influence of majority opinion has long been used to encourage changes in attitudes and behaviors. In the traditional human-human domain, studies have shown their effectiveness in contexts such as health interventions (Skalski and Tamborini, 2007; Zhang et al., 2015) and sustainability efforts (Athanasiadis and Mitkas, 2005; Vossen et al., 2009). In human-computer interaction, persuasive systems have been designed across numerous domains like health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021), and social welfare (Shi et al., 2020), aiming to improve user behavior or change their attitudes through persuasion and informational education. Yet, these implementations are limited to single-agent interactions. The potential of multi-agent systems that collectively effect changes in user behavior or opinion remains unexplored.
@@
The potential of AI agents in enacting persuasive effects is further supported by research in the HCI domain, which has highlighted the significant social influence individual AI agents exert, shaping users’ attitudes (Jakesch et al., 2023) and behaviors (Shi et al., 2020; Zhang et al., 2020). However, such research has focused on the design characteristics of individual agents, such as how their perceived identity as a chatbot or human impacts outcomes (Shi et al., 2020). Little attention has been given to how the number of agents affects the effectiveness of these systems. In contrast, HCI research on multi-agent system research has largely revolved around enhancing task performance (Guo et al., 2024; Chen et al., 2023b), either by helping users complete tasks more effectively or improving the technical performance of the system itself (Du et al., 2023; Chen et al., 2023a). The indirect effects of social influence caused by these multi-agent systems has been largely overlooked. The impact of social influence is less relevant to systems where the primary goal is to improve either the clarity of the agent’s role in the discussion or the technical performance of the agents themselves. In scenarios where the system needs to both convey information and persuade a user of a certain viewpoint, the effect of multiple vs. single agents is critically underexplored.
@@
Therefore, given 1) the increasing prevalence of multi-agent systems and interactions with multiple social bots online, 2) the known social influence that AI agents can exert, and 3) the established impacts of group social influence among humans that may be mirrored in human-AI interactions, there is 4) significant potential for both positive and negative effects caused by the collective social influence created by a group of AI agents. In this study, we address this issue by asking: Can interacting with multiple virtual agents create social influence that shifts our opinions? Specifically, we explore whether interactions with multiple agents can lead to stronger opinion changes, induce varied social influence, and identify demographic factors that increase susceptibility to AI influence, aiming to inform ethical guidelines for multi-agent systems.
@@
To investigate these questions, we conducted an experiment (n = 94) where participants discussed their stances on two social topics with one, three, or five agents. Each participant was randomly assigned to one condition, with discussion content kept consistent across groups, differing only in the number of agents presenting it. Agents presented arguments that aligned with the participant’s stance on one topic and opposed it on the other. Survey data and open-ended responses were collected before and after each conversation round.
@@
Through quantitative and qualitative analysis, we found that the multi-agent setup significantly influenced opinion shifts. Participants showed stronger opinion changes toward the agents when they disagreed and greater polarization when agents agreed with them. However, the shift did not intensify as the number of agents increased from 3 to 5. Instead, the 5-agent group led to a sense of polarization, with more participants rejecting the agents’ arguments and moving further from the agents’ positions. Additionally, multi-agent groups heightened perceptions of social influence, particularly normative pressure to align with the majority. Open-ended responses indicated that participants expressed a desire for affiliation within the group, which contributed to social influence and opinion shifts in the 3- and 5-agent conditions. Demographic analysis further revealed that younger participants were more susceptible to multi-agent influence.
@@
Our study makes the following contributions to future research in the HCI and CSCW communities:
@@
  • We enhance the understanding of the differences in social influence between single-agent and multi-agent systems. By observing participants’ feelings of social exclusion and their desire to engage, we identified an implicit sense of affiliation with a group of agents, expanding existing social influence theories to include human-agent interactions.
  • We provide design suggestions for using multi-agent setups to more effectively convey persuasive strategies. By leveraging the normative pressure exerted by multiple agents, future systems could foster emotional engagement in participants, promoting more lasting and internalized changes in attitudes and behaviors.
@@ 2. Related Work
2.1. Multi-agent Interaction
AI agents are increasingly integrated in our daily life. These advanced AI systems are capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). Recent research and industry applications have leveraged the conversational and reasoning capabilities of LLMs to create problem-solving tools (Wang et al., 2024a; Roy et al., 2024; Ning et al., 2024; Yu et al., 2024; Guan et al., 2024), educational assistants (Lieb and Goel, 2024), and novel search interfaces (Ma et al., 2024; Yang et al., 2025). In commercial settings, many companies deploy AI agents on social media to assist with branding and customer service (Xu et al., 2017; Hu et al., 2018; Jiang et al., 2022), while others embed agents in their applications for assisting writing (Dhillon et al., 2024), brainstorming (Rezwana and Maher, 2023), or gaming (Sidji et al., 2024).
Despite this, there are many tasks that single LLM-based agents struggle to handle effectively. Individual LLMs can lack sufficient domain expertise to handle specialized tasks (Ge et al., 2023), and may have insufficiently robust reasoning capabilities. Inspired by existing psychological theories (Smoliar, 1991), studies have investigated techniques such as ”debating” (Du et al., 2023) or ”reconciling” (Chen et al., 2023a) between multiple agents, to enhance reasoning by leveraging diverse perspectives. This has driven growing interest in ”multi-agent collaboration,” where agents work together to improve task performance (Guo et al., 2024; Chen et al., 2023b). Fields such as natural language processing (Guo et al., 2024), software engineering (Cardoso and Ferrando, 2021), and robotics (Chen et al., 2024) have leveraged the collective intelligence of multiple agents to simulate a group dynamic and improve task performance (Qian et al., 2023; Hong et al., 2023; Xiong et al., 2023). Multi-agent setups have also been used to simulate different human roles, akin to multiple workers interacting (Park et al., 2022, 2023b) or collaborating on a task (Light et al., 2023; Aher et al., 2023).
@@ 2.1. Multi-agent Interaction
Given the potential prevalence of human and multi-agent collaboration in everyday life, the interaction between multiple agents and humans remains underexplored in the HCI and CSCW domains. Previous work has a limited focus on multi-agent systems in HCI, being largely designed to improve user outcomes - either by reducing cognitive load in information gathering (Jiang et al., 2023) or to provide information from diverse sources to improve decision making (Tan and Liew, 2022; Park et al., 2023a; Beinema et al., 2021; Chaves and Gerosa, 2018). These studies mostly highlight opportunities or challenges of multi-agent interfaces, such as users feeling overwhelmed or confused by multiple agents providing different inputs (Chaves and Gerosa, 2018; Tan and Liew, 2022; Park et al., 2023a). In contrast, no existing work to our knowledge has investigated if and how a group of agents affects users through social influence mechanisms. Such effects are distinct from the cognitive and decision-making effects in previous studies, yet are potentially as important.
In this study, we investigate the scenario: what happens if multiple agents are perceived as social actors, and they all share the same opinion on a topic, just like a group of people might? This situation can arise, for instance, when companies deploy multiple agents on social media with a unified stance. While the influence of a single opinionated bot is well-documented (Jakesch et al., 2023; Shi et al., 2020; Zhang et al., 2020), the impact of being surrounded by multiple like-minded bots is less understood. Considering the growing prevalence of multi-agent systems, it is crucial to investigate how they might shape people’s opinions. Drawing from existing theories on human behavior within social groups, we hypothesize several potential effects, which will be discussed in the next section.
@@ 2.2. Social Influence Theory
When multiple people express the same opinion, others in the group often feel compelled to agree. This phenomenon is known as ”social influence” (Cialdini and Goldstein, 2004), which comprises the ways in which individuals adjust their behavior to meet the demands of a social environment. This theory is well-known and widely studied in social psychology, and has been applied across different contexts, such as marketing (Salganik et al., 2006; Zhu et al., 2012; Teo et al., 2019), health intervention (Skalski and Tamborini, 2007; Zhang et al., 2015), sustainability (Athanasiadis and Mitkas, 2005; Vossen et al., 2009), and political discussions (Price et al., 2006).
@@
Social influence is a broad term that manifests in various forms. Two main types of social influence are normative influence, the desire to ”conform with the expectations of another”, and informational influence, the desire to ”accept information obtained from another as evidence about reality” (Deutsch and Gerard, 1955). Normative influence relates to many common types of social influence, including conformity (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), socialization (Richters and Waters, 1991; Oostveen et al., 1996), and social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011). More recent literature has termed this mechanism affiliation (Cialdini and Goldstein, 2004), or a desire to fit in with the majority: assimilation into a social group is often desired, and agreeing with the majority opinion can expedite this process. Naturally, normative social influence is more common among individuals who form a collective social group (Deutsch and Gerard, 1955), which hence creates an environment in which one might want to be a part of. Informational influence, on the other hand, relates to the concept of accuracy (Cialdini and Goldstein, 2004), or the natural instinct to gravitate towards objectively correct opinions as expressed by a group. This is distinct from normative influence as individuals may ”accept an opponent’s beliefs as evidence about reality even though one has no [other] motivation to agree with him” (Deutsch and Gerard, 1955). Despite this, social influence can still persuade individuals to accept an opinion, as the perceived consensus (Cialdini and Goldstein, 2004) of a larger number of people expressing the same viewpoint may suggest that that viewpoint is accurate.
A key motivation for this study is that it is unclear whether these two forms of influence also apply to human-AI interactions, as AI agents possess different characteristics from humans. With informational influence, a key distinction between human-human and human-AI interactions is that in human groups, each person represents a distinct source of knowledge and possible truth. Conversely, with AI agents, it is trivial to set up multiple agents based on the same source of data, meaning that the number of agents presenting an opinion has no relation to their objective accuracy. AI agents may have broader but more generic knowledge, especially in factual or domain-specific areas, yet lack personal or context-specific insights. Regarding normative influence, AI agents typically lack key social attributes and characteristics (Ozdemir et al., 2023) such as empathy, credibility, and authority, which are typically foundational to social influence in human groups (Cialdini and Goldstein, 2004). Consequently, it remains uncertain whether multiple AI agents can evoke the same sense of normative pressure or emotional affiliation that human groups do. Nevertheless, the well-established tendency of humans to perceive such agents as distinct social actors (Nass et al., 1994; Li et al., 2025a) raises the possibility of both informational and normative influence from a group of agents.
In this study, we investigate whether a group of AI agents can create a similar sense of a social group that implicitly encourages a user to join it by way of agreeing with their opinions.
To understand how social influence operates in human-human contexts—specifically, the conditions under which it is strengthened or weakened—prior research has primarily focused on two dimensions: the source of influence (i.e., characteristics of social groups that amplify or reduce influence) and the recipient (i.e., individual traits that make one more or less susceptible to influence). On the source side, studies have examined various group-level factors that impact persuasive power, such as group dynamics (Affinity et al., 2003) and group identity (Spears, 2021), and group size (Wilder, 1977). Among these, group dynamics—particularly the presence of group consensus—has been shown to play a crucial role in strengthening both normative and informational social influence (Gardikiotis et al., 2005; Affinity et al., 2003). When a group presents a unified stance, individuals are more likely to conform either because they wish to fit in (normative influence) or because they infer that the group holds valid information (informational influence) (Affinity et al., 2003). Drawing on these findings, we designed our multi-agent system to simulate a group with a consensus opinion, representing the most direct way to induce social influence in a controlled setting.
On the recipient side, individual attributes—such as gender (Carli, 2001), age (Steinberg and Monahan, 2007), and education level (Abu-Shanab, 2011)—have been shown to significantly shape susceptibility to social influence and their trust in technology. For example, recent work by Hackenburg et al. (Hackenburg and Margetts, 2024) found that age, gender, and education level influence how individuals respond to persuasive messages generated by LLMs. Understanding which demographic factors are associated with greater susceptibility is essential for the design of responsible AI systems, in order to prevent undue influence or potential harm to specific populations (Hackenburg and Margetts, 2024). Building on these insights, our study investigates whether users’ demographic characteristics—such as age, gender, and education—affect their susceptibility to social influence in the context of interacting with multiple AI agents.
@@
In HCI research, a few studies (e.g. (Adam et al., 2021)) have examined the social influence that individual chatbots can exert on humans. In this study, we aim to conduct, to our knowledge, the first investigation of the possibility of both normative and informational social influence caused by groups of AI agents.
2.3. Conversational Agent in Online Discussion